# Machine Learning-Based Detection of Pump-and-Dump Schemes in Real-Time

### v1

@@
Abstract
Cryptocurrency markets often face manipulation through prevalent pump-and-dump (P&D) schemes, where self-organized Telegram groups, some exceeding two million members, artificially inflate target cryptocurrency prices. These groups sell premium access to inside information, worsening information asymmetry and financial risks for subscribers and all investors. This paper presents a real-time prediction pipeline to forecast target coins and alert investors to possible P&D schemes. In a Poloniex case study, the model accurately identified the target coin among the top five from 50 random coins in 24 out of 43 (55.81%) P&D events. The pipeline uses advanced natural language processing (NLP) to classify Telegram messages, identifying 2,079 past pump events and detecting new ones in real-time. Our analysis also evaluates the susceptibility of token standards—ERC-20, ERC-721, BRC-20, Inscriptions, and Runes—to manipulation and identifies exchanges commonly involved in P&D schemes.
Index Terms:
@@ Abstract
I Introduction
Blockchain is an immutable, decentralized ledger that eliminates the need for a central authority, enabling secure and transparent transactions. It facilitates the continuous tracking of ownership through digital tokens, which act as units of value. Tokens can store value, serve as participation rewards, provide voting rights, or enable interaction with products and services [1, 2, 3]. However, tokens are also susceptible to market manipulations, making their attributes critical to this study.
Cryptocurrency markets promise high returns and decentralized control but remain vulnerable to market manipulations like pump-and-dump (P&D) schemes. These schemes, orchestrated via messaging platforms such as Telegram and Discord, artificially inflate a cryptocurrency’s price, enabling organizers to profit at the expense of unsuspecting investors. A recent report by Chainalysis [4] revealed that 24% of tokens launched in 2022 exhibited characteristics typical of P&D schemes, underlining the systemic risks these activities pose to market integrity and the urgent need for effective detection and prevention mechanisms.
@@ I Introduction
Although traditional analysis of trading volume and token price can provide valuable indicators for the emergence of P&D schemes [5], recent advances in large language models (LLMs) [6, 7, 8] allow for better detection and understanding of these manipulations. LLMs enable detailed analysis of Telegram messages, often the primary medium through which P&D organizers communicate with participants. By extracting patterns and identifying key linguistic cues in these messages, LLMs can complement conventional market data analysis and offer a more comprehensive detection approach.
@@
The introduction of new token standards, such as Ordinals [9], Inscriptions [10], and Runes [11, 12, 13], on Bitcoin, further complicates the landscape of P&D schemes. These standards spread to the EVM-chains [14], raising questions about the susceptibility of tokens beyond the widely targeted ERC-20 standard on Ethereum and other EVM-compatible chains. While ERC-20 tokens have historically been prime targets due to their widespread adoption and ease of trading, understanding whether and how emerging standards are exploited is crucial for developing holistic anti-manipulation strategies.
@@
Early studies on P&D schemes focused on identifying price and volume anomalies after pumps occurred [5], achieving moderate success but underscoring the need for more sophisticated models. Real-time detection approaches, such as anomaly detection using machine learning [15], have shown promise but suffer from latency issues, as seen in the 30-minute lag of certain methods [16]. More recent methods [17] leverage Random Forest and AdaBoost classifiers to detect anomalies within seconds of pump initiation.
Target coin prediction aims to identify manipulated coins before a pump. Xu et al. [18] developed Random Forest models to predict pump likelihood using market metrics, while Hu et al. [19] introduced sequence-based deep-learning models leveraging channel-specific features. Despite advancements, existing approaches largely rely on historical data and lack integration of high-frequency order book data, limiting their real-time applicability.
@@
  3. 3.
    Evaluating the susceptibility of emerging token standards like BRC-20 and Runes to P&D schemes.
@@
The paper is structured as follows: Section II introduces the relevant token standards. Section III presents the pipeline system architecture for real-time prediction of P&D schemes, Section IV details the data collection process. Empirical results are presented in Section V, followed by a discussion in Section VI, and conclusions in Section VII.
II Background
To provide context for this research, we begin with a concise introduction to various token standards.
II-1 ERC-20
ERC-20 is a widely adopted standard for fungible tokens on Ethereum, enabling interoperability across decentralized applications (dApps). Tokens under this standard can represent anything from assets to utility tokens, facilitating broad use cases [20, 21]. BEP-20, a similar standard on Binance Smart Chain, offers lower transaction fees and congestion, enhancing its appeal [22].
II-2 ERC-721
ERC-721 defines a standard for non-fungible tokens (NFTs), which are unique and indivisible. NFTs enable the representation of ownership for digital assets like art and collectibles, ensuring their provenance and authenticity through blockchain’s immutable records [23, 24].
II-3 Ordinals and Inscriptions
Ordinals enable individual satoshis to carry unique identifiers, granting them subjective value based on collector sentiment [9]. Inscriptions enhance this concept by attaching data like images or text to satoshis, transforming them into unique digital artifacts that can be traded [25].
II-4 BRC-20
The BRC-20 standard introduces fungible tokens on Bitcoin, leveraging the Ordinals protocol to inscribe metadata onto satoshis. While BRC-20 enables token creation and transfer, its programmability is limited compared to Ethereum-based standards [10]. This simplicity highlights its experimental nature, relying on off-chain tools for effective tracking and management.
II-5 Runes
Runes embed fungible tokens directly into Bitcoin’s UTXO system, allowing for efficient token creation, minting, and transfer. The protocol ensures integrity through a burning mechanism for invalid transactions and maintains a smaller on-chain footprint, addressing scalability concerns [11, 12, 13].
By exploring these token standards, this research provides insights into the attributes that make certain tokens more vulnerable to P&D schemes.
III Pipeline System Architecture
Figure 1: Architecture of the Real-time P&D Detection Pipeline.
@@ III Pipeline System Architecture
IV Data Collection and Methodology
TABLE I: Summary of data sources utilized in P&D detection pipeline, including purpose, scope, and relevant details.
@@ IV Data Collection and Methodology
The pipeline incorporates 365,982 rows of daily OHLCV (open, high, low, close, volume) price data, social metrics, and market capitalization data for 4,643 cryptocurrencies. Additionally, it stores between 5 and 44 million rows of order book metrics and between 2 and 5 million rows of trade metrics per exchange. These metrics, crucial for target coin prediction, are calculated and updated every few seconds. Table I presents a comprehensive overview of these data sources, which are elaborated upon in this section. Data was collected from 30 days up to 1 hour before each pump event.
IV-A Pump Message Collection
Previous studies on target coin prediction [18] primarily relied on data from the now-defunct PumpOlymp [^1] PumpOlymp was previously active under https://pumpolymp.com/ website to identify pump channels on Telegram. In the absence of a comparable successor, we utilized TGStats [26] to search for specific keywords within channel titles and descriptions, including crypto, pump, dump, P&D, and signal, along with their combinations. Identified channels were manually validated for active promotion of P&D schemes, resulting in the identification of 43 active channels. From these, we extracted complete message histories, creating a dataset currently comprising 91,295 messages. Telegram was chosen as the primary focus, as previous studies have established its role as the main platform for P&D activities [18].
IV-B Pump Message Labeling
Manual examination of the collected messages revealed six distinct categories of P&D-related messages, aligned with the pump anatomy described in Section I. The dataset was preprocessed to normalize characters, spaces, and emojis. User-tags were replaced with the generic token @USER, and non-informative URLs were replaced with UNKNOWN_URL. A list of prominent centralized exchanges (CEXs), decentralized exchanges (DEXs), and defunct platforms such as HotBit and Cryptopia was compiled to tag messages mentioning exchanges with special tokens (_CEX or _DEX).
To automate P&D detection, we labeled approximately 25% (21,092 messages) of the dataset across the six categories using the GPT-4o model via the OpenAI API. Detailed prompts and parameter settings are available in Appendix A-A. Manual verification and re-labeling were performed to enhance label accuracy, focusing on critical classes: Pump Announcement, Countdown, Target Coin Release, and Delay/Cancellation. Only a small number of messages required re-labeling, demonstrating the reliability of GPT-4o as a baseline model.
TABLE II: Distribution of Labels Assigned to Telegram Messages.
@@ IV-B Pump Message Labeling
The labeled data was used to fine-tune a BERTweet [27] model. Table II summarizes the label distribution after cleaning. To address label imbalance, a weighted loss function was applied during training. The dataset was split into 80% for training and 20% for validation and testing. Duplicate messages, commonly reused across channels (e.g., ”1 HOUR UNTIL THE PUMP…”), were excluded from the test set to ensure unbiased evaluation metrics. The performance of the model is included in Appendix in Table VI.
@@
IV-C Market Data
We analyze 365,982 rows of daily OHLCV data for 4,643 cryptocurrencies, including price, trading volume, and market capitalization metrics critical for identifying P&D targets. Data is collected via API calls to exchanges such as KuCoin and MEXC, supplemented by the CCXT library [28], and automated through a daily cron job. Preprocessed data is stored in a database and used in filter models for predictive analysis.
Challenges in historical data accuracy arose due to non-unique coin symbols and missing data for 403 of the 1,045 unique symbols, often due to delistings or defunct exchanges. To address this, our real-time pipeline incorporates self-reported market capitalization data from CoinMarketCap [29]. Additionally, derivative tokens, such as leveraged and inverse tokens, were excluded using a pattern-matching approach, ensuring that only primary cryptocurrencies were considered.
IV-D Social Data
Social sentiment significantly impacts cryptocurrency prices [30]. Using the LunarCrush API [31], we collect metrics such as likes, shares, and social dominance. Features such as rolling interaction means and social dominance ratios are engineered to explore correlations with market trends.
IV-E Order Book Data
@@ IV-F Trade Data
IV-G Target Coin Normalization
@@
Figure 2: Relative Z-Scores with Pumped Coins Highlighted
@@
We tested the model at time offsets of 20 seconds, 40 seconds, and 1 minute before the pump. As shown in Table III, TOP5 and TOP10 performances declined with increasing time offsets, highlighting the importance of near-real-time predictions. When using both trade and order book data, the TOP5 performance decreased from 55.81% (20 seconds) to 19.05% (1 minute). Similarly, using only trade data, TOP5 performance dropped from 46.51% to 9.52%.
TABLE III: TOP5 and TOP10 Performance with Different Time Offsets
@@
V Empirical Results
First, the model labeled messages and extracted details such as pump times, exchanges, and target coins. This yielded a dataset of 2,079 P&D events between December 2017 and September 2024. As shown in Table VIII, most events occurred on centralized exchanges (CEX), with Hotbit leading despite ceasing operations in May 2023. Recent activity has shifted to LATOKEN, XT, and Poloniex, as illustrated in Figure 3.
Figure 3: Temporal Distribution of P&D Events by Exchange.
V-A Quantification of P&D Schemes and Target Coins
Figure 4: Market Caps, Price, and Volume Distribution of Pumped Coins per Channel
We analyzed market cap and price data from CoinCodex [32] for pumped coins. Supplementary Figure 4 in Appendix highlights that most pumped coins belong to the small- and mid-cap categories. Many coins were targeted multiple times, with some becoming frequent targets. Table VII in Appendix lists the five most pumped coins, led by TOKKI, which was pumped nine times across four exchanges by five different channels.
V-A1 Volume
Our analysis reveals a significant impact of pump events on the trade volumes of targeted coins. Supplementary Figures 13 and 13 in Appendix illustrate the average daily trade volumes for approximately 700 pump events on KuCoin and Poloniex over an eight-day period, with the eighth day representing the pump event. The substantial increase in trade volume on the pump day compared to a typical trading day is strikingly evident.
@@ V-A1 Volume
We further analyzed average trade volumes grouped by the Telegram channels orchestrating these pump events. Supplementary Figures 13 and 13 in Appendix highlight the variation in trade volume increases across different channels on the pump day. Interestingly, the observed variations cannot be solely attributed to easily measurable factors such as the number of subscribers per channel. This indicates that additional, less apparent factors play a role in determining the magnitude of a pump event’s impact.
V-A2 Price Spikes

### v2

# Real-Time Machine Learning Detection of Telegram-Based Pump-and-Dump Schemes

@@
Abstract.
Cryptocurrency markets often face manipulation through prevalent pump-and-dump (P&D) schemes, where self-organized Telegram groups, some exceeding two million members, artificially inflate target cryptocurrency prices. These groups sell premium access to inside information, worsening information asymmetry and financial risks for subscribers and all investors. This paper presents a real-time prediction pipeline to forecast target coins and alert investors to possible P&D schemes. In a Poloniex case study, the model accurately identified the target coin among the top five from 50 random coins in 24 out of 43 (55.81%) P&D events. The pipeline uses advanced natural language processing (NLP) to classify Telegram messages, identifying 2,079 past pump events and detecting new ones in real-time.
@@ Abstract.
$c$s: Computing methodologies Machine learningccs: Applied computing Economicsconference: ACM CCS Workshop on Decentralized Finance and Security; October 17, 2025; Taipei, Taiwanbooktitle: Proceedings of the ACM CCS Workshop on Decentralized Finance and Security (DeFi’25), October 17, 2025, Taipei, Taiwan
1. Introduction
Blockchain is an immutable, decentralized ledger that eliminates the need for a central authority, enabling secure and transparent transactions. It facilitates the continuous tracking of ownership through digital tokens, which act as units of value. Tokens can store value, serve as participation rewards, provide voting rights, or enable interaction with products and services (Pilkington, 2016; Freni et al., 2020; Gogol et al., 2024). However, tokens are also susceptible to market manipulations, making their attributes critical to this study.
Cryptocurrency markets promise high returns and decentralized control but remain vulnerable to market manipulations like pump-and-dump (P&D) schemes. These schemes, orchestrated via messaging platforms such as Telegram and Discord, artificially inflate a cryptocurrency’s price, enabling organizers to profit at the expense of unsuspecting investors. A recent report by Chainalysis (Chainalysis, 2023) revealed that 24% of tokens launched in 2022 exhibited characteristics typical of P&D schemes, underlining the systemic risks these activities pose to market integrity and the urgent need for effective detection and prevention mechanisms.
@@ 1. Introduction
Although traditional analysis of trading volume and token price can provide valuable indicators for the emergence of P&D schemes (Kamps and Kleinberg, 2018), recent advances in large language models (LLMs) (Kalyan, 2024; Chang, Yupeng and Wang, Xu and Wang, Jindong and Wu, Yuan and Yang, Linyi and Zhu, Kaijie and Chen, Hao and Yi, Xiaoyuan and Wang, Cunxiang and Wang, Yidong and Ye, Wei and Zhang, Yue and Chang, Yi and Yu, Philip S. and Yang, Qiang and Xie, Xing, 2024; Wang et al., 2024) allow for better detection and understanding of these manipulations. LLMs enable detailed analysis of Telegram messages, often the primary medium through which P&D organizers communicate with participants. By extracting patterns and identifying key linguistic cues in these messages, LLMs can complement conventional market data analysis and offer a more comprehensive detection approach.
@@
The introduction of new token standards, such as Ordinals (Rodarmor, 2023b), Inscriptions (Li et al., 2024), and Runes (Rodarmor, 2023c; Vermaak, 2023; KuCoin, 2024), on Bitcoin, further complicates the landscape of P&D schemes. These standards spread to the EVM-chains (Messias et al., 2024), raising questions about the susceptibility of tokens beyond the widely targeted ERC-20 standard on Ethereum and other EVM-compatible chains. While ERC-20 tokens have historically been prime targets due to their widespread adoption and ease of trading, understanding whether and how emerging standards are exploited is crucial for developing holistic anti-manipulation strategies.
@@
Early studies on P&D schemes focused on identifying price and volume anomalies after pumps occurred (Kamps and Kleinberg, 2018), achieving moderate success but underscoring the need for more sophisticated models. Real-time detection approaches, such as anomaly detection using machine learning (Mansourifar et al., 2020), have shown promise but suffer from latency issues, as seen in the 30-minute lag of certain methods (Victor and Hagemann, 2019). More recent methods (La Morgia et al., 2023) leverage Random Forest and AdaBoost classifiers to detect anomalies within seconds of pump initiation.
Target coin prediction aims to identify manipulated coins before a pump. Xu et al. (Xu and Livshits, 2019) developed Random Forest models to predict pump likelihood using market metrics, while Hu et al. (Hu et al., 2023) introduced sequence-based deep-learning models leveraging channel-specific features. Despite advancements, existing approaches largely rely on historical data and lack integration of high-frequency order book data, limiting their real-time applicability.
@@
  3. (3)
@@
The paper is structured as follows: Section 2 introduces the relevant token standards. Section 3 presents the pipeline system architecture for real-time prediction of P&D schemes, Section 4 details the data collection process. Empirical results are presented in Section 5, followed by a discussion in Section 6, and conclusions in Section 7.
2. Preliminaries
To provide context for this research, we briefly introduce main token standards.
ERC-20. ERC-20 is a widely adopted standard for fungible tokens on Ethereum, enabling interoperability across decentralized applications (dApps). Tokens under this standard can represent anything from assets to utility tokens, facilitating broad use cases (Morales et al., 2023; Vogelsteller and Buterin, 2015). BEP-20, a similar standard on Binance Smart Chain, offers lower transaction fees and congestion, enhancing its appeal (AlphaPoint, 2024).
ERC-721. ERC-721 defines a standard for non-fungible tokens (NFTs), which are unique and indivisible. NFTs enable the representation of ownership for digital assets like art and collectibles, ensuring their provenance and authenticity through blockchain’s immutable records (Barrington and Merrill, 2022; Entriken et al., 2018).
Ordinals and Inscriptions. Ordinals enable individual satoshis to carry unique identifiers, granting them subjective value based on collector sentiment (Rodarmor, 2023b). Inscriptions enhance this concept by attaching data like images or text to satoshis, transforming them into unique digital artifacts that can be traded (Rodarmor, 2023a).
BRC-20. The BRC-20 standard introduces fungible tokens on Bitcoin, leveraging the Ordinals protocol to inscribe metadata onto satoshis. While BRC-20 enables token creation and transfer, its programmability is limited compared to Ethereum-based standards (Li et al., 2024). This simplicity highlights its experimental nature, relying on off-chain tools for effective tracking and management.
Runes. Runes embed fungible tokens directly into Bitcoin’s UTXO system, allowing for efficient token creation, minting, and transfer. The protocol ensures integrity through a burning mechanism for invalid transactions and maintains a smaller on-chain footprint, addressing scalability concerns (Rodarmor, 2023c; Vermaak, 2023; KuCoin, 2024).
3. Pipeline System Architecture
Figure 1. Architecture of the Real-time P&D Detection Pipeline.
@@ 3. Pipeline System Architecture
4. Data Collection and Methodology
Table 1. Summary of data sources utilized in P&D detection pipeline, including purpose, scope, and relevant details.
@@ 4. Data Collection and Methodology
The pipeline incorporates 365,982 rows of daily OHLCV (open, high, low, close, volume) price data, social metrics, and market capitalization data for 4,643 cryptocurrencies. Additionally, it stores between 5 and 44 million rows of order book metrics and between 2 and 5 million rows of trade metrics per exchange. These metrics, crucial for target coin prediction, are calculated and updated every few seconds. Table 1 presents a comprehensive overview of these data sources, which are elaborated upon in this section. Data was collected from 30 days up to 1 hour before each pump event.
4.1. Pump Message Collection
Previous studies on target coin prediction (Xu and Livshits, 2019) primarily relied on data from the now-defunct PumpOlymp [^1] PumpOlymp was previously active under https://pumpolymp.com/ website to identify pump channels on Telegram. In the absence of a comparable successor, we utilized TGStats (TGStat, 2024) to search for specific keywords within channel titles and descriptions, including crypto, pump, dump, P&D, and signal, along with their combinations. Identified channels were manually validated for active promotion of P&D schemes, resulting in the identification of 43 active channels. From these, we extracted complete message histories, creating a dataset currently comprising 91,295 messages. Telegram was chosen as the primary focus, as previous studies have established its role as the main platform for P&D activities (Xu and Livshits, 2019).
4.2. Pump Message Labeling
Manual examination of the collected messages revealed six distinct categories of P&D-related messages, aligned with the pump anatomy described in Section 1. The dataset was preprocessed to normalize characters, spaces, and emojis. User-tags were replaced with the generic token @USER, and non-informative URLs were replaced with UNKNOWN_URL. A list of prominent centralized exchanges (CEXs), decentralized exchanges (DEXs), and defunct platforms such as HotBit and Cryptopia was compiled to tag messages mentioning exchanges with special tokens (_CEX or _DEX).
To automate P&D detection, we labeled approximately 25% (21,092 messages) of the dataset across the six categories using the GPT-4o model via the OpenAI API. Detailed prompts and parameter settings are available in Appendix A.1. Manual verification and re-labeling were performed to enhance label accuracy, focusing on critical classes: Pump Announcement, Countdown, Target Coin Release, and Delay/Cancellation. Only a small number of messages required re-labeling, demonstrating the reliability of GPT-4o as a baseline model.
Table 2. Distribution of labels assigned to Telegram messages.
@@ 4.2. Pump Message Labeling
Table 3. PumpBERT classification performance.
Metric Value
Overall Metrics
F1 weighted average 0.982
Precision 0.982
Recall 0.982
Label-specific F1 Scores
Pump Announcement 0.976
Countdown 0.993
Target Coin Release 0.995
Pump Results 0.970
Delay/Cancellation 1.000
Noise 0.971
The labeled data was used to fine-tune a BERTweet (Nguyen et al., 2020) model. Table 3 summarizes the label distribution after cleaning. To address label imbalance, a weighted loss function was applied during training. The dataset was split into 80% for training and 20% for validation and testing. Duplicate messages, commonly reused across channels (e.g., ”1 HOUR UNTIL THE PUMP…”), were excluded from the test set to ensure unbiased evaluation metrics. The performance of the model is included in Table 3.
@@
4.3. Market Data
We analyze 365,982 rows of daily OHLCV data for 4,643 cryptocurrencies, including price, trading volume, and market capitalization metrics critical for identifying P&D targets. Data is collected via API calls to exchanges such as KuCoin and MEXC, supplemented by the CCXT library (Kroitor and contributors, nd), and automated through a daily cron job. Preprocessed data is stored in a database and used in filter models for predictive analysis.
Challenges in historical data accuracy arose due to non-unique coin symbols and missing data for 403 of the 1,045 unique symbols, often due to delistings or defunct exchanges. To address this, our real-time pipeline incorporates self-reported market capitalization data from CoinMarketCap (Coi, 2024). Additionally, derivative tokens, such as leveraged and inverse tokens, were excluded using a pattern-matching approach, ensuring that only primary cryptocurrencies were considered.
4.4. Social Data
Social sentiment significantly impacts cryptocurrency prices (Wołk, 2019). Using the LunarCrush API (Inc., 2024), we collect metrics such as likes, shares, and social dominance. Features such as rolling interaction means and social dominance ratios are engineered to explore correlations with market trends.
4.5. Order Book Data
@@ 4.6. Trade Data
4.7. Target Coin Normalization
@@
Figure 2. Relative Z-Scores with Pumped Coins Highlighted
@@
We tested the model at time offsets of 20 seconds, 40 seconds, and 1 minute before the pump. As shown in Table 4, TOP5 and TOP10 performances declined with increasing time offsets, highlighting the importance of near-real-time predictions. When using both trade and order book data, the TOP5 performance decreased from 55.81% (20 seconds) to 19.05% (1 minute). Similarly, using only trade data, TOP5 performance dropped from 46.51% to 9.52%.
Table 4. TOP5 and TOP10 Performance with Different Time Offsets
@@
5. Empirical Results
First, the model labeled messages and extracted details such as pump times, exchanges, and target coins. This yielded a dataset of 2,079 P&D events between December 2017 and September 2024. As shown in Table 10 in Appendix, most events occurred on centralized exchanges (CEX), with Hotbit leading despite ceasing operations in May 2023. Recent activity has shifted to LATOKEN, XT, and Poloniex, as illustrated in Figure 3.
Figure 3. Temporal Distribution of P&D Events by Exchange.
5.1. Quantification of P&D Schemes and Target Coins
Figure 4. Market Caps, Price, and Volume Distribution of Pumped Coins per Channel
We analyzed market cap and price data from CoinCodex (CoinCodex, 2024) for pumped coins. Supplementary Figure 4 in Appendix highlights that most pumped coins belong to the small- and mid-cap categories. Many coins were targeted multiple times, with some becoming frequent targets. Table 11 in Appendix lists the five most pumped coins, led by TOKKI, which was pumped nine times across four exchanges by five different channels.
5.1.1. Volume
Our analysis reveals a significant impact of pump events on the trade volumes of targeted coins. Supplementary Figures 18 and 18 in Appendix illustrate the average daily trade volumes for approximately 700 pump events on KuCoin and Poloniex over an eight-day period, with the eighth day representing the pump event. The substantial increase in trade volume on the pump day compared to a typical trading day is strikingly evident.
@@ 5.1.1. Volume
We further analyzed average trade volumes grouped by the Telegram channels orchestrating these pump events. Supplementary Figures 18 and 18 in Appendix highlight the variation in trade volume increases across different channels on the pump day. Interestingly, the observed variations cannot be solely attributed to easily measurable factors such as the number of subscribers per channel. This indicates that additional, less apparent factors play a role in determining the magnitude of a pump event’s impact.
5.1.2. Price Spikes